OpenStudy (anonymous):

PLEASE help...me... √2 cosxtanx−tanx=0 Solve for 0

OpenStudy (anonymous):

It would be easier if you made everything sin and cos. sqrt{2} cos(x)*(sin(x)/cos(x))-(sin(x)/cos(x))=0 simplifying, it becomes sqrt(2)sin(x)-(sin(x)/cos(x))=0 because there is a \[\sqrt{2}\] in the first part, one of the answers has to be pi/4. that way, the \[\sqrt{2}\] in front of sin(x) and the \[\sqrt{2}\] in the denominator of sin(pi/4) cancels out. the other answer has to be 3pi/4 because of the exact same reason